OpenStudy (anonymous):

|-15| + 32 = 50y - |-37| - 49y is it -13??

OpenStudy (mertsj):

\[15+32=50y-37-49y\]

OpenStudy (mertsj):

\[47=y-37\]

OpenStudy (mertsj):

\[84=y\]

OpenStudy (anonymous):

|-15| + 3^2 = 50y - |-37| - 49y

OpenStudy (mertsj):

How is it you can post 3^2 now but couldn't in the original posting?

OpenStudy (anonymous):

I corrected it.. sorry it was a typo

OpenStudy (mertsj):

y=61